# Generating Probabilistic Scenario Programs from Natural Language

### v2

@@ Abstract
For cyber-physical systems (CPS), including robotics and autonomous vehicles, mass deployment has been hindered by fatal errors that occur when operating in rare events. To replicate rare events such as vehicle crashes, many companies have created logging systems and employed crash reconstruction experts to meticulously recreate these valuable events in simulation. However, in these methods, ”what if” questions are not easily formulated and answered. We present ScenarioNL, an AI System for creating scenario programs from natural language. Specifically, we generate these programs from police crash reports. Reports normally contain uncertainty about the exact details of the incidents which we represent through a Probabilistic Programming Language (PPL), Scenic. By using Scenic, we can clearly and concisely represent uncertainty and variation over CPS behaviors, properties, and interactions. We demonstrate how commonplace prompting techniques with the best Large Language Models (LLM) are incapable of reasoning about probabilistic scenario programs and generating code for low-resource languages such as Scenic. Our system is comprised of several LLMs chained together with several kinds of prompting strategies, a compiler, and a simulator. We evaluate our system on publicly available autonomous vehicle crash reports in California from the last five years and share insights into how we generate code that is both semantically meaningful and syntactically correct.
@@ 1 Introduction
Figure 1: Architecture diagram of the ScenarioNL Compound AI System.
@@
Robust validation of the safety and performance of end-to-end autonomous driving systems in simulation requires an extensive and diverse dataset of simulations modeling more dangerous and unusual driving scenarios. This allows for more precise measures of vehicle behavior and safety in higher-risk scenarios that may not appear as frequently in on-road testing or log-based simulations. Within this domain, there exists a plethora of detailed driving data available through datasets such as nuScenes (Caesar et al., 2020) or Berkeley Deep Drive 100K (Yu et al., 2020). There are also video-only datasets specifically detailing accidents such as the Dashcam Accident Dataset (Chan et al., 2017) and Car Crash Dataset (Bao et al., 2020). Recent works such as that by Adewopo et al. (2024) integrate some of these existing datasets into a large video dataset from multiple source such as dashcams, sensors, traffic surveillance, and small clips from selected video streaming platforms.
@@
Despite this abundance of publicly-available data, only a fraction contains the corresponding sensor data, and an even smaller portion is relevant to the autonomous driving domain (traffic surveillance video may not help an autonomous vehicle to be a better driver). Moreover, the high-risk situations most important to validate for autonomous driving systems often appear the least frequently in real-world recorded data. These scenarios include rare events and unsafe or unusual driving behaviors by surrounding vehicles, which may result in collisions or dangerous scenarios. These underrepresented scenarios are important for testing not only to train systems to avoid them, but they are also necessary for ethical research into decision making when a system must act when a collision is unavoidable (Robinson et al., 2021). While, datasets such as Car Crash (Bao et al., 2020) and Dashcam Accident (Chan et al., 2017) have collected these higher-risk scenarios they are only videos without sensor data and are limited in size and duration. For example, the Car Crash Dataset only has 1500 examples that are 2 seconds long thus missing the critical events leading up to a crash. To the best of our knowledge, there is no dataset containing full driving scenarios in which the recorded vehicles were involved in an accident or in a challenging situation. Thus, one promising but challenging opportunity is to generate a realistic dataset containing dangerous driving scenarios.
With current methods, Kalra & Paddock estimated that autonomous vehicles (AV) would have to log over 11 billion miles to beat the human driver error rate. This would take hundreds of years for today’s fleets to log. Consequently, autonomous driving companies have continuously invested in simulation-based testing of AVs over the last few years instead of purely on-road testing. Notably, Waymo has invested in more than one simulator for years (Hawkins, 2021) and has recently released yet another, WayMax (Gulino et al., 2023), built from their Open Motion Dataset. Aside from the tractability issues, testing of specific components or end-to-end autonomous driving systems in simulation mitigates the risks of human harm or system damage present in real-world testing.
@@
Given the critical need for simulation based testing and the lack of data pertaining to the rare events needed to gain confidence in our autonomous driving systems, we ask how can we get this data? We observe that within the United States, regulators offer a dataset of a different form than that of sensor or video recordings. For example, the California Department of Motor Vehicles mandates that all AVs must report all crash incidents of AVs. The National Transportation Safety Board (NTSB) and the Nevada State Legislature also provide reports for crashes of significant monetary value. Thus we ask, how can we leverage the available reports to reconstruct dangerous and unusual automotive situations in simulation? This question was asked by Scanlon et al. (2021). The authors collected all Waymo crashes in Arizona over 10 years and paid a team of ”trained experts” to perform exact crash reconstruction. Their dataset of 72 crashes were then used to perform counterfactual ”what-if” simulations to learn insights about these rare events. Similar to Waymo, we reconstruct crash scenes in simulation.
Our Compound Artificial Intelligence (AI) System (Khattab et al., ) called, ScenarioNL, uses Scenic (Fremont et al., 2019; 2023) to represent driving scenarios as Probabilistic Programs where uncertainty in crash details are represented through probability distributions. The choice of Scenic as the target language is best illustrated with an example.
@@
Figure 2: Example ScenarioNL output and its respective CARLA simulator visualization. The natural language description was parsed from a California DMV Crash report. It specifies the collision between a cyclist and paused autonomous vehicle. After the collision, the cyclist leaves the scene without exchanging information.
@@ 1.1 Scenic as a domain specific language
Figure 2 illustrates the Scenic program generalizing a collision between a vehicle and a bicyclist. Scenic itself is a Python-like language augmented with several language constructs to make representing such scenarios easier. For example, Scenic includes several constructs for specifying that an object is visible, on a particular orient point, or relative to another object / point. Additionally, one can declaritively specify distributions – with all objects modeled by Scenic, e.g. cars, bicycles, weather, etc, coming with prior distributions. For example, the distribution of car models and colors attempts to match known vehicle statistics. Finally, Scenic enables declaring the behavior of other road users via a co-routine like API.
All together, Scenic enables succinctly capturing common automotive scenarios. This has several benefits. First, the resulting programs are often human readable and editable. Second, because the programs are succinct, they are more likely to fit into the context of a language model. Finally, because Scenic has a formal semantics, we can guide existing LLMs to generate Scenic programs by using domain specific compiler feedback.
@@ 1.2 Contributions
To scale the reproduction of Scenic programs, ScenarioNL uses compositional prompting techniques, a compiler, and Large Language Models (LLM) to develop these programs in a low resource setting such as Scenic.
@@
    Our system, ScenarioNL, is a Compound AI System (gupta_shift_nodate) leveraging LLMs for generating Scenic Probabilistic Programs from natural language text. Unlike Scanlon et al. (2021), our system is fully automated and does not require human experts. It can handle a variety of natural language from simple text descriptions to full PDF crash reports thus surpassing the capabilities of existing diffusion-based approaches such that by Zhong et al. (2023). Furthermore, our system does is not restricted by template based strategies such as those by Deng et al. (2023) and Wang et al. (2023b). It can theoretically represent any scenario. Furthermore, with additional natural language descriptions of AVs, we can represent any scenario and not just those with crashes.
@@
    Qualitative Analysis of our system reveals the difficulty of generating Domain Specific Languages (DSL) for LLMs. We share insights on how composing multiple strategies can alleviate some of the difficulties. Our methods are not unique to Scenic, and do not assume the existence of labeled data thus enabling the opportunity for other DSL generation.
@@
    We further augment our approach by taking a systems’ point-of-view. We incorporate feedback from existing tools such as the Scenic Compiler to further guide LLMs towards semantically and syntactically correct Scenic programs. Our system consistently outputs syntactically-correct programs at rates of up to $90\%$.
@@ 2 Related Work
2.1 Code Generation with Large Language Models
@@
Most of these model’s are fine-tuned variants of a foundational model that was trained on a wider body of text such as Common Crawl. For instance Codex is ”a GPT language model fine-tuned on publicly available code from GitHub” and it’s authors, ”study its Python code-writing capabilities” (Chen et al., 2021). The authors used 159GB of Python files. Such a vast amount of data is not available for many Domain Specific Languages (DSL) such as Lean (Moura & Ullrich, 2021), and Isabelle (Jaskelioff & Merz, 2005) (the only DSL to have proved the correctness of a full operating system). In these cases, fine-tuning is not an option as the number of examples is ever so small as compared to the amount used in prior works. In the case of Isabelle, there are less than 1000 archived proofs (noa, ). Thus, new techniques are required. One such work is LeanDojo (Yang et al., 2023b). The authors collect nearly 100k proof tactic examples, and train a small transformer from scratch to output individual proof tactics at every state of a proof. While their method works for an interactive theorem prover, it is not able to write a full proof on its own. Furthermore, on an even less commonly used DSL, such as Scenic, there are less than 30 publicly available code examples thus making both fine-tuning and the LeanDojo approach (a ReProver model) infeasible.
3 ScenarioNL
@@
While works such as Kolla et al. (2022) have demonstrated that exact crash reconstruction is possible, Scanlon et al. (2021) showed that it is when we ask ”what-if” questions that we gain insights on how the accident occurred and how we can avoid it. Scenic (Fremont et al., 2019), a Scenario-based Probabilistic Programming Language (PPL), strikes a balance between the two. With a concise syntax, Scenic allows one to represent the spatial-temporal relationships between objects such as cars on a map. Uncertainty about the environment and object properties can be encoded as probability distributions thus allowing for a natural way to represent uncertainty and ”what-if” questions. Scenic also allows for program evaluation with realistic physics by interfacing with simulators. In our setting, we choose CARLA (Dosovitskiy et al., 2017), a vehicle simulator illustrated in our introductory example, Figure 2. By representing scenarios as Scenic programs and using ScenarioNL to generate these programs, we can create hundreds of scenarios answering thousands of ”what-if” questions.
@@
Problem setup. For simplicity, we assume that the inputs and outputs of our system will belong to a common alphabet $\Sigma$, e.g., tokenized ASCII symbols. We say that a string $P\in\Sigma^{*}$ is a Scenic program if it conforms to the formal grammar/syntax defined in Fremont et al. (2019). We say that a Scenic program $P$ is valid if it produces no runtime errors. Let SCENIC denote the set of all valid Scenic programs. The goal of ScenarioNL is to sample valid Scenic programs, $P\in\text{SCENIC}$, given a natural language prompt $x\in\Sigma^{*}$.
@@
We note two features about this problem: First, because Scenic is a domain specific language well suited to modeling automotive scenarios, the code generation needed is often shorter than a general purpose language. On the other hand, because Scenic is not very common in internet scale datasets, there is an apriori low probability that an LLM will generate valid Scenic.
@@ 3.1 System Architecture
ScenarioNL generates Scenic programs from a collection of publicly available crash reports involving AVs. By running a single command, the user can specify any supported LLM as a backbone and compositional prompting strategies. The modular system design supports integration of any new LLM or prompting strategy with the addition of a single new adapter file. The system automatically caches any intermediate outputs to conserve resources and API calls. An architecture diagram is provided in Figure 1
@@ 3.2 Input: Autonomous Vehicle Collision Reports
The system input is a web-scraped collection of over 500 collision reports involving AVs over the last five years, hosted by the California Department of Motor Vehicles (DMV, 2023). Scenarios detailed in the autonomous vehicle collision reports may involve interactions with other vehicles, pedestrians, cyclists, or other objects. ScenarioNL uses optical character recognition (OCR) to extract any relevant data from the collision report PDFs. The final natural language input contains collision information including weather and environmental conditions, sustained vehicle damage, and description of the events leading up to the accident. The input data represents a vast range of scenarios with potentially multiple cars, pedestrians, and collisions. Next, due to the wide variety of scenarios possible, we manually classify the reports:
@@
    Easy Reports that contain one dynamic agent (ex. an AV was rear-ended by another agent while the AV was legally at rest) on a public road.
@@
    Medium There are two dynamic agents (ex. an AV and cyclist), or the scenario is in a road configuration not supported in a default CARLA map (ex. a parking lot).
@@
    Hard There are three or more dynamic agents, scenarios involving aggressive drivers, or scenarios involving advanced driving maneuvers (ex. illegal vehicle overtaking).
@@ 3.3 Prompting
Modular LLM Prompting ScenarioNL accepts the natural language scene description as input in order to formulate a prompt given a user-specified approach and language model. The modular LLM prompting mechanism allows for ease of evaluation and code generation on a wide range of approaches. This may range from unique decoding constraints to few-shot examples retrieved from a vector database. ScenarioNL allows users to specify a model and prompting technique in the one-line gen_scenic command. The full list of strategies supported are described in detail in the Appendix.
Compositional LLM Prompting Some prompting techniques are better for different reasons at different times. For example, few-shot prompting can be used to each an LLM a specific output format whereas, chain-of-thought (Wei et al., 2023) can be used to elicit better reasoning capabilities from an LLM. Our key insight is that we should separate reasoning from coding as much as possible and leverage constrained decoding to help guide the LLMs as much as possible. We implement the following composition of prompts:
@@
    ToT We again return to our experts to debate about the answers of the prior questions. Each question is answered directly where possible (ex. weather condition or road structure such as 4 way intersection) or in the form of a probability distribution, where confidence may be represented as variance (ex. Vehicle speed can be represented as a Gaussian distribution with mean and variance).
@@
    Few-Shot & Constrained Decoding Given that our experts were able to reason about the relevant objects and uncertainty in the original description, we now construct the Scenic program by parts (constants/parameters, agent behaviors, and spatiotemporal relations). These program parts are further fed into the next parts of the ScenarioNL system. The constrained decoding is implemented through LMQL (Beurer-Kellner et al., 2023) — a package that allows users to express constraints as python code rather than having to directly manipulate token logit biases. We use LMQL version 0.7 with gpt-3.5-turbo-instruct, the most recent OpenAI model compatible with LMQL custom constraints. Finally, we leverage Retrevial Augmented Generation (RAG) & Hypothetical Document Embeddings (HyDE) in each section to further help the LLM. We refer the reader to the appendix for a deeper explanation on RAG and HyDE.

### v3

# ScenicNL: Generating Probabilistic Scenario Programs from Natural Language

@@ Abstract
For cyber-physical systems, including robotics and autonomous vehicles, mass deployment has been hindered by fatal errors that occur when operating in rare events. To better understand failure modes, companies meticulously recreate rare crash events in simulation, but current methods do not easily allow for exploring ”what if” scenarios which could reveal how accidents might have been avoided. We present ScenicNL, an AI system that generates probabilistic scenario programs from natural language. Given the abundance of documented failures of autonomous vehicles due to regulatory requirements, we apply ScenicNL to police crash reports, providing a data-driven approach to capturing and understanding these failures. By using a probabilistic language such as Scenic, we can clearly and concisely represent such scenarios of interest and easily ask “what if” questions. We demonstrate how commonplace prompting techniques with Large Language Models are incapable of generating code for low-resource languages such as Scenic. We propose an AI system via the composition of several prompting techniques to extract the reasoning abilities needed to model probability distributions around the uncertainty in the crash events. Our system then uses Constrained Decoding and tools such as a compiler and simulator to produce scenario programs in this low-resource setting. We evaluate our system on publicly available autonomous vehicle crash reports in California from the last five years and share insights into how we generate code that is both semantically meaningful and syntactically correct. Finally, we release our code and a collection of over 500 crash reports from the California Department of Motor Vehicles.
@@ 1 Introduction
Robust verification and validation of the safety and performance of autonomous artificial intelligence (AI) enabled cyber-physical systems (CPS) requires an integrated approach spanning formal modeling, simulation, and scalable verification and testing methods (Seshia et al., 2022). Simulation-based verification requires performing an extensive and diverse set of simulations including scenarios that are more dangerous and unusual than typical scenarios. As an example, for autonomous driving systems, this allows for evaluation of safety in higher-risk scenarios that may not appear as frequently in on-road testing or log-based simulations. Within this domain, there exists a plethora of detailed driving data available through datasets such as nuScenes (Caesar et al., 2020) or Berkeley Deep Drive 100K (Yu et al., 2020). Recent works such as that by Adewopo et al. (2024) integrate some of these existing datasets into a large video dataset from multiple sources such as dashcams, sensors, traffic surveillance, and small clips from selected video streaming platforms.
@@
Despite the abundance of publicly-available data, only a fraction contains the corresponding sensor data, and an even smaller portion is relevant to the autonomous driving domain (e.g. traffic surveillance video may not help an autonomous vehicle to be a better driver). Moreover, the high-risk situations most important to validate for autonomous driving systems often appear the least frequently in real-world recorded data. These scenarios include rare events, and unsafe or unusual driving behaviors by surrounding vehicles, which may result in collisions or dangerous scenarios. These underrepresented scenarios are important for testing not only to train systems to avoid them, but they are also necessary for ethical research into decision making when a system must act when a collision is unavoidable (Robinson et al., 2021). While datasets such as Car Crash (Bao et al., 2020) and Dashcam Accident (Chan et al., 2017) have collected these higher-risk scenarios, they are only videos without sensor data and are limited in size and duration. For example, the Car Crash Dataset has 1500 examples that are 2 seconds long thus missing the critical events leading up to a crash. To the best of our knowledge, there is no dataset containing full driving scenarios in which the recorded vehicles were involved in an accident or in a challenging situation. Thus, one promising but challenging opportunity is to generate a realistic dataset containing dangerous driving scenarios.
Figure 1: Architecture diagram of the ScenicNL compound AI system.
@@
Given the critical need for simulation based testing and the lack of data pertaining to the rare events needed to gain confidence in our autonomous driving systems, we ask how can we get this data? We observe that within the United States, regulators offer a dataset of a different form than that of sensor or video recordings. For example, the California Department of Motor Vehicles mandates that all AVs must report all crash incidents. The National Transportation Safety Board and the Nevada State Legislature also provide reports for crashes of significant monetary value. Thus we ask, how can we leverage available reports to reconstruct dangerous and unusual automotive situations in simulation? This question was asked by Scanlon et al. (2021). The authors collected all Waymo crashes in Arizona over 10 years and paid a team of ”trained experts” to perform exact crash reconstruction. Their dataset of 72 crashes were then used to perform counterfactual ”what-if” simulations to learn insights about these rare events. Similar to Waymo, we reconstruct crash scenes in simulation.
@@
Figure 2: Example ScenicNL output and its respective CARLA simulator visualization. The natural language description was parsed from a California DMV Crash report. It specifies the collision between a cyclist and paused autonomous vehicle.
In this paper, we address the gap of a missing dataset that contains dangerous and rare driving scenarios by introducing ScenicNL, a system capable of automatically generating such a dataset. Our system generates probabilistic programs representing these scenarios of interest from natural language descriptions. By generating probabilistic programs, our system can account for uncertainty and variability in the descriptions by modeling events and properties as probability distributions i.e. a single program can represent many variations of the same scenario. These scenario programs when sampled yield concrete instantiations in a driving simulator. Unlike prior works, we perform this reconstruction in a completely automated fashion. Moreover, we are able to generate programs for a wide variety of scenarios by foregoing coding templates that past works have used.
In order to produce these scenario programs without templates or human experts, our system leverages Large Language Models (LLMs). The challenge in developing ScenicNL is thus, how we can get LLMs to write code in Scenic, a Domain Specific Language (DSL). We demonstrate how commonly used prompting techniques are insufficient to write code in our DSL. Without relying on an extensive knowledge base (vector storage) or fine-tuning, we develop a compound artificial intelligence (AI) system (Zaharia et al., 2024) and share general insights on how the composition of multiple prompting strategies and the integration of tools can help tackle the problem of LLMs writing code for any DSL. The choice of Scenic as the target language is best illustrated with an example.
@@ 1.1 Scenic as a domain specific language
Figure 2 illustrates the Scenic program generalizing a collision between a vehicle and a bicyclist. Scenic itself is a Python-like language augmented with several language constructs to make representing such scenarios easier. For example, Scenic includes several constructs for specifying that an object is visible, on a particular orient point, or relative to another object / point. Additionally, one can declaritively specify distributions – with all objects modeled by Scenic, e.g. cars, bicycles, weather, etc., coming with prior distributions. For example, the distribution of car models and colors attempts to match known vehicle statistics. Finally, Scenic enables declaring the behavior of other road users via a co-routine like API.
All together, Scenic enables succinctly capturing common automotive scenarios. This has several benefits. First, the resulting programs are often human readable and editable. Second, because the programs are succinct, they are more likely to fit into the context of a language model. Third, Scenic is an open-source system that is well-supported by verification and testing tools such as VerifAI Dreossi et al. (2019). Finally, because Scenic has a formal semantics, we can guide existing LLMs to generate Scenic programs by using domain specific compiler feedback.
@@ 1.2 Contributions
To scale the reproduction of Scenic programs, ScenicNL uses compositional prompting techniques, a compiler, and Large Language Models (LLM) to develop these programs in a low resource setting such as Scenic.
@@
    Our system, ScenicNL, is a Compound AI System (Zaharia et al., 2024) leveraging LLMs for generating Scenic Probabilistic Programs from natural language text. Unlike Scanlon et al. (2021), our system is fully automated and does not require human experts. It can handle a variety of natural language from simple text descriptions to full PDF crash reports thus surpassing the capabilities of existing diffusion-based approaches such that by Zhong et al. (2023). Furthermore, our system does is not restricted by template based strategies such as those by Deng et al. (2023) and Wang et al. (2023b). It can theoretically represent any scenario. With additional natural language descriptions of AVs, we can represent any scenario and not just those with crashes.
@@
    Qualitative Analysis of our system reveals the difficulty of generating DSLs for LLMs. We share insights on how composing multiple strategies can alleviate some of the difficulties. Our methods are not unique to Scenic, and do not assume the existence of labeled data thus enabling the opportunity for other DSL generation.
@@
    We further augment our approach by taking a systems’ approach. We incorporate feedback from existing tools such as the Scenic Compiler to further guide LLMs towards semantically and syntactically correct Scenic programs. Our system consistently outputs syntactically-correct programs at rates of up to $90\%$.
{mdframed}
[nobreak=true] While our contributions are demonstrated on the autonomous driving domain, we emphasize that Scenic itself is a language applicable to any AI-enabled CPS or robotics domain, and our ScenicNL system is also independent of the specific application we demonstrate our results on.
@@ 2 Related Work
Most of these model’s are fine-tuned variants of a foundational model that was trained on a wider body of text such as Common Crawl. For instance Codex is ”a GPT language model fine-tuned on publicly available code from GitHub” and it’s authors, ”study its Python code-writing capabilities” (Chen et al., 2021). The authors used 159GB of Python files. Such a vast amount of data is not available for many DSL such as Lean (Moura & Ullrich, 2021), and Isabelle (Jaskelioff & Merz, 2005). In these cases, fine-tuning is not an option as the number of examples is ever so small as compared to the amount used in prior works. In the case of Isabelle, there are less than 1000 archived proofs (noa, ). Thus, new techniques are required. One such work is LeanDojo (Yang et al., 2023b). The authors collect nearly 100k proof tactic examples, and train a small transformer from scratch to output individual proof tactics at every state of a proof. While their method works for an interactive theorem prover, it is not able to write a full proof on its own. Furthermore, on an even less commonly used DSL, such as Scenic, there are less than 30 publicly available code examples thus making both fine-tuning and the LeanDojo approach (a ReProver model) infeasible.
3 ScenicNL
@@
While works such as Kolla et al. (2022) have demonstrated that exact crash reconstruction is possible, Scanlon et al. (2021) showed that it is when we ask ”what-if” questions that we gain insights on how the accident occurred and how we can avoid it. Scenic (Fremont et al., 2019), a Probabilistic Programming Language (PPL), strikes a balance between the two. With a concise syntax, Scenic allows one to represent the spatial-temporal relationships between objects such as cars on a map. Uncertainty about the environment and object properties can be encoded as probability distributions thus allowing for a natural way to represent uncertainty and ”what-if” questions. Scenic also allows for program evaluation with realistic physics by interfacing with simulators. In our setting, we choose CARLA (Dosovitskiy et al., 2017), a vehicle simulator illustrated in our introductory example, Figure 2. By representing scenarios as Scenic programs and using ScenicNL to generate these programs, we can create hundreds of scenarios answering thousands of ”what-if” questions.
@@
Problem setup. For simplicity, we assume that the inputs and outputs of our system will belong to a common alphabet $\Sigma$, e.g., tokenized ASCII symbols. We say that a string $P\in\Sigma^{*}$ is a Scenic program if it conforms to the formal grammar/syntax defined in Fremont et al. (2019). We say that a Scenic program $P$ is valid if it produces no runtime errors. Let SCENIC denote the set of all valid Scenic programs. The goal of ScenicNL is to sample valid Scenic programs, $P\in\text{SCENIC}$, given a natural language prompt $x\in\Sigma^{*}$.
@@
We note two features about this problem: First, because Scenic is a DSL well suited to modeling scenarios for autonomous systems and robotics, the code generation needed is often shorter than a general purpose language. On the other hand, because Scenic is not very common in internet scale datasets, there is an apriori low probability that an LLM will generate valid Scenic.
@@ 3.1 System Architecture
ScenicNL generates Scenic programs from a collection of publicly available crash reports involving AVs. By running a single command, the user can specify any supported LLM as a backbone and compositional prompting strategies. The modular system design supports integration of any new LLM or prompting strategy with the addition of a single new adapter file. The system automatically caches any intermediate outputs to conserve resources and API calls. An architecture diagram is provided in Figure 1
@@ 3.2 Input: Autonomous Vehicle Collision Reports
The system input is a web-scraped collection of over 500 collision reports involving AVs over the last five years, hosted by the California Department of Motor Vehicles (DMV, 2023). Scenarios detailed in the autonomous vehicle collision reports may involve interactions with other vehicles, pedestrians, cyclists, or other objects. ScenicNL uses optical character recognition (OCR) to extract any relevant data from the collision report PDFs. The final natural language input contains collision information including weather and environmental conditions, sustained vehicle damage, and description of the events leading up to the accident. The input data represents a vast range of scenarios with potentially multiple cars, pedestrians, several modes of locomotion (e.g. skateboarding), and collisions. Next, due to the wide variety of scenarios possible, we manually classify the reports:
@@
    Easy Reports that contain one dynamic agent (e.g. an AV was rear-ended by another agent while the AV was legally at rest) on a public road.
@@
    Medium There are two dynamic agents (e.g. an AV and cyclist), or the scenario is in a road configuration not supported in a default CARLA map (e.g. a parking lot).
@@
    Hard There are three or more dynamic agents, scenarios involving aggressive drivers, or scenarios involving advanced driving maneuvers (e.g. illegal vehicle overtaking).
@@ 3.3 Prompting
Figure 3: Overview of ScenicNL’s Compositional Prompting. Blue rectangles utilize Tree-of-Thought with Role Playing and a single example program. Red hexagrams use Constrained Decoding to generate correct Scenic code. A single-shot program is generated from partially correct code, reasoning in natural language, and program subparts which were corrected (if necessary) with a compiler in the loop. A Few-Shot full program is then generated and looped with a compiler to emit a final program.
Modular LLM Prompting ScenicNL accepts the natural language scene description as input in order to formulate a prompt given a user-specified approach and language model. The modular LLM prompting mechanism allows for ease of evaluation and code generation on a wide range of approaches. This may range from unique decoding constraints to few-shot examples retrieved from a vector database. ScenicNL allows users to specify a model and prompting technique in the one-line gen_scenic command. The full list of strategies supported are described in detail in the Appendix.
Compositional LLM Prompting Some prompting techniques are better for different reasons at different times. For example, few-shot prompting can be used to demonstrate a specific output format whereas, chain-of-thought (Wei et al., 2023) can be used to elicit better reasoning capabilities from an LLM. Our key insight is that we should separate reasoning from coding as much as possible. When working in the natural language space, we can employ multiple personas to further elicit world knowledge from a single LLM thus better identifying and describing uncertainties in a scenario. Furthermore, we ask these personas to generate their own questions about uncertainly in the crash report descriptions which they then literately attempt to answer. By using multiple personas and self-generated questions, we find that the quality and quantity of identified uncertainties from a report increases over a single LLM query. We also observe that leveraging constrained decoding to help guide the LLMs to syntactically correct code where possible further improves system performance as we don’t have to hope for correct code which must be validated by a compiler downstream. We implement the following composition of prompts:
@@
    ToT We again return to our experts to debate about the answers of the prior questions. Each question is answered directly where possible (e.g. weather condition or road structure such as 4 way intersection) or in the form of a probability distribution, where confidence may be represented as variance (e.g. Vehicle speed can be represented as a Gaussian distribution with mean and variance).
@@
    Few-Shot & Constrained Decoding Given that our experts were able to reason about the relevant objects and uncertainty in the original description, we now construct the Scenic program by parts (constants/parameters, agent behaviors, and spatiotemporal relations). These program parts are further fed into the next parts of the ScenicNL system. The constrained decoding is implemented through LMQL (Beurer-Kellner et al., 2023) — a package that allows users to express constraints as python code rather than having to directly manipulate token logit biases. We use LMQL version 0.7 with gpt-3.5-turbo-instruct, the most recent OpenAI model compatible with LMQL custom constraints. Finally, we leverage Retrevial Augmented Generation (RAG) & Hypothetical Document Embeddings (HyDE) in each section to further help the LLM. We refer the reader to the appendix for a deeper explanation on RAG and HyDE.
An overview of our prompts is shown in Figure 3. The full prompts and questions from the figure are provided in the appendix.